Matthew K. Gray

Work Address MIT Room E15-022 20 Ames Street Cambridge, MA 02139

(617) 253–2383 Home Address 64 Webster Ave Cambridge, MA 02141 (617)-497–3934

Experience Personal Information Architecture Group, MIT Media Lab

Cambridge, MA

May, 1997 – present Research into networking infrastructure for “Things That Think”, a large scale project on ubiquitous computing in everyday objects. Research responsibilities include those described below, as well as supervising undergraduates, team management, and high level project design.

Physics and Media Group, MIT Media Lab Cambridge, MA April, 1996 – May, 1997 Research into electrostatic intrabody signaling. Research responsibilities include hardware and software design, parts acquisition, implementation, testing, presentation, and handling sponsor relationships.

net.Genesis Corporation Cambridge, MA April, 1994 – January, 1996 Founder and Chief Technologist for a company offering server side software tools for web developers. Developed initial versions of products, did product design and strategic planning. Other responsibilities included coding, sales support, marketing support, customer service, and management of engineering teams.

O’Reilly and Associates Camgridge, MA May, 1994 – August, 1994 Consulting work focusing on producing advertiser reports and performance enhancements to the GNN web server.

MIT Media Lab Cambridge, MA January, 1994 – June, 1994 Developing software for a project to provide customized newspapers to MIT students from Associated Press and other news sources. Coding primarily in C and perl.

MIT Planetary Science Department (EAPS) Cambridge, MA May, 1993 – August, 1993 Observation and data analysis of light curves from Koronis Asteroids. Familiarity with observational precedures and astronomical image processing.

MIT Radio Astronomy Group Cambridge, MA January, 1993 – May, 1993 Worked on signal processing of light curves attempting to measure time delays in gravitational lense images. Work included C programming and work with image processing packages. Continuation of earlier work.

Brookhaven National Lab Brookhaven, NY May, 1992 – August, 1992 Data analysis and presentation for particle physics work, specifically, seeking a rare kaon decay (E787 at BNL). Developed Software in C and Fortran. Continuation of earlier work.

Education Massachussetts Institute of Technology Media Laboratory

Cambridge, MA

M.S. in Media Technology, May 1999

Massachussetts Institute of Technology Cambridge, MA B.S. in Physics, May 1997

Publications Build a Web Site, Prima Publishing, June, 1995. A book published while with net.Genesis covering technical and design aspects of managing web sites.

“Growth of the Web Report,”, Web (http://www.mit.edu/people/mkgray/net), continually updated. A report first published in June, 1993 detailing historical growth in the number of web sites, based on original research. Extensively cited, including Newsweek, PC Week, PC world, Communications Week, and many others.

“Web Watch,”, Internet World, 1995. A monthly column written by me and others at net.Genesis.

Organizations, Proficiencies, Interests

SIPB - Student Information Processing Board January, 1992 – present Member. A volunteer organization providing computing services to the MIT community. I was Chairman in 1993. I set up one of world’s first 100 web servers in May 1993 as part of this group.

Apache Group Member. Volunteer development group of the Apache Web Server, the most popular web server in the world.

Computer Languages Exceptional proficiency: C, perl, Java. Proficiency: C++, sh, VB, Fortran, tcl/tk, many others.

Interests Basic conversational French speaker, Ballroom dance, Recreational computing, Volunteer activities.